Consider an investment that pays $41.49 in year 1, and then stabilizes and pays $4.81 every year forever after that (the first cash flow is in year 2) This firm does not intend to grow and has an interest rate (required rate of return) of 5%. What is the present value of this investment opportunity? Give your answer to two decimals
Answer #1

horizon value = perpetual payment/rate of return = 4.81/0.05=96.2

Value in year 1 = horizon value+Cash flow in year 1 = 41.49+96.2=137.69

Future value = present value*(1+ rate)^time
137.69 = Present value*(1+0.05)^1
Present value = 131.133
Answer #2

Given:

  • Cash flow in Year 1 (CF₁): $41.49

  • Perpetual cash flow from Year 2 onward (CF): $4.81 per year

  • Discount rate (r): 5% or 0.05


Step 1: Present Value of Year 1 Cash Flow

The present value of the cash flow received in Year 1 is simply its discounted value:

PVYear 1=CF1(1+r)=41.491.0539.51

Step 2: Present Value of the Perpetuity (Starting in Year 2)

The perpetuity formula gives the value one year before the first cash flow, which is Year 1 in this case (since the perpetuity starts in Year 2):

PVPerpetuity at Year 1=CFr=4.810.05=96.20

Now, discount this value back to the present (Year 0):

PVPerpetuity at Year 0=96.20(1+r)=96.201.0591.62

Step 3: Total Present Value

Add the present values of the two components:

PVTotal=PVYear 1+PVPerpetuity at Year 0=39.51+91.62=131.13

Answer:

The present value of the investment opportunity is 131.13


answered by: anonymous